OpenStudy (anonymous):

Compute the sum 4(6)+6(-8)+(-8)(-10).

OpenStudy (camerondoherty):

4 times 6 = a 6 times -8 = b -8 times -10 = c a + b + c = ?

OpenStudy (camerondoherty):

Do you get what im saying?

OpenStudy (anonymous):

just wait I'm figuring it out

OpenStudy (camerondoherty):

Ok how about this: what is 4 times 6?

OpenStudy (anonymous):

24

OpenStudy (camerondoherty):

what is six times -8

OpenStudy (anonymous):

is the answer for the problem 56

OpenStudy (camerondoherty):

yes